# Markov kernels in Mathlib’s probability library

### v1

@@ 6 Information theory and divergences
In this last section on applications of kernels to the formalization other probabilistic notions, we will discuss information-theoretic quantities: entropy and the Kullback-Leibler divergence (KL). Those examples are not from Mathlib, but from other projects (although the definition of KL is now in Mathlib, klDiv). They should be seen as experiments on how the kernel approach can be applied to other notions, and not as a definitive design choices.
In the PFR project [And+23], we formalized the entropy of a probability measure $\mu$ on a finite set $\mathcal{X}$ as $H(\mu)=-\sum_{x\in\mathcal{X}}\mu(\{x\})\log\mu(\{x\})$, and the conditional entropy of a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ with respect to a probability measure $\mu$ on $\mathcal{X}$ as
@@
but we also proved that the conditional entropy of $X$ given $Y$ is equal to $H_{k}\left(P_{X|Y}\mid P_{Y}\right)$ and derived many properties for the kernel version of entropy, before then specializing to the conditional entropy. For example, the chain rule for kernel entropy can be stated as
@@
The Kullback-Leibler divergence definition now in Mathlib was developed in the TestingLowerBounds project [LD24]. A highlight of the project is a proof of the data-processing inequality for f-divergences, a class of divergences between probability distributions that includes KL, as well as the chain rule for KL. The Kullback-Leibler divergence between two probability measures $\mu,\nu$ on a measurable space $\mathcal{X}$ is defined as
@@
if $\mu$ is absolutely continuous with respect to $\nu$, and $\infty$ otherwise. The usual formulation of the chain rule for KL states that if $\mu,\nu$ are two probability measures on $\mathcal{X}$ and $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ are two Markov kernels, then
@@
$\mathrm{KL}(\mu\otimes_{m}\kappa,\nu\otimes_{m}\eta)$ $=\mathrm{KL}(\mu,\nu)+KL(\mu\otimes_{m}\kappa,\mu\otimes_{m}\eta)\>.$
@@
For the Rényi divergence $R_{\alpha}(\mu,\nu)=\frac{1}{\alpha-1}\log\left(\int_{x}\left(\frac{d\mu}{d(\mu+\nu)}(x)\right)^{\alpha}\left(\frac{d\nu}{d(\mu+\nu)}(x)\right)^{1-\alpha}d(\mu+\nu)\right)$ for $\alpha\in(0,1)$, the composition-product and integral definitions do not coincide. It is remarkable that the conditional definition that enjoys a form of the chain rule is not the integral definition, but the composition-product of kernels version $R_{\alpha}(\mu\otimes_{m}\kappa,\nu\otimes_{m}\eta)$ [PW25, Chapter 7]. We thus expect Markov kernels to play an important role in the formalization of probability divergences in Mathlib.

### v2

@@ 6 Information theory and divergences
In this last section on applications of kernels to the formalization of other probabilistic notions, we will discuss information-theoretic quantities: entropy and the Kullback-Leibler divergence (KL). Those examples are not from Mathlib, but from contributions of the author to other projects (although the definition of KL is now in Mathlib, klDiv). They should be seen as experiments on how the kernel approach can be applied to other notions, and not as definitive design choices.
In the PFR project [Anderson_Formalization_of_the_2023], we formalized the entropy of a probability measure $\mu$ on a finite set $\mathcal{X}$ as $H(\mu)=-\sum_{x\in\mathcal{X}}\mu(\{x\})\log\mu(\{x\})$, and the conditional entropy of a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ with respect to a probability measure $\mu$ on $\mathcal{X}$ as
@@
but we also proved that the conditional entropy of $X$ given $Y$ is equal to $H_{k}\left(P_{X|Y}\mid P_{Y}\right)$ (recall that $P_{X|Y}$ is a conditional distribution, see Section 3.2) and derived many properties for the kernel version of entropy, before then specializing to the conditional entropy. For example, the chain rule for kernel entropy can be stated as
@@
The Kullback-Leibler divergence definition now in Mathlib was developed in the TestingLowerBounds project [TestingLowerBounds]. A highlight of the project is a proof of the data-processing inequality for f-divergences, a class of divergences between probability distributions that includes KL, as well as the chain rule for KL. The Kullback-Leibler divergence between two probability measures $\mu,\nu$ on a measurable space $\mathcal{X}$ is defined as
@@
if $\mu$ is absolutely continuous with respect to $\nu$, and $\infty$ otherwise. The usual formulation of the chain rule for $\mathrm{KL}$ states that if $\mu,\nu$ are two probability measures on $\mathcal{X}$ and $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ are two Markov kernels, then
@@
$\mathrm{KL}(\mu\otimes_{m}\kappa,\nu\otimes_{m}\eta)$ $=\mathrm{KL}(\mu,\nu)+\mathrm{KL}(\mu\otimes_{m}\kappa,\mu\otimes_{m}\eta)\>.$
@@
For the Rényi divergence $R_{\alpha}(\mu,\nu)=\frac{1}{\alpha-1}\log\left(\int_{x}\left(\frac{d\mu}{d(\mu+\nu)}(x)\right)^{\alpha}\left(\frac{d\nu}{d(\mu+\nu)}(x)\right)^{1-\alpha}d(\mu+\nu)\right)$ for $\alpha\in(0,1)$, the composition-product and integral definitions do not coincide. It is remarkable that the conditional definition that enjoys a form of the chain rule is not the integral definition, but the composition-product of kernels version $R_{\alpha}(\mu\otimes_{m}\kappa,\nu\otimes_{m}\eta)$ [polyanskiy2025information, Chapter 7]. We thus expect Markov kernels to play an important role in the formalization of probability divergences in Mathlib.
7 Related work
We discuss formalizations of transition kernels in other proof assistants, as well as an application of kernels in Mathlib to prove the Ionescu-Tulcea theorem and define the product measure of an infinite family of probability measures.
7.1 Markov kernels in other proof assistants
The formalization of Markov kernels in other proof assistants is mostly motivated by the semantics of probabilistic programming languages [staton2020probabilistic], while our goal is to provide a general-purpose theory of kernels for probability theory. In Isabelle/HOL, [hirata2023semantic] formalized quasi-Borel spaces and defined Markov kernels, finite and s-finite kernels. In Rocq/Coq, [affeldt2025semantics] also formalized Markov, finite and s-finite kernels, and additionally $\sigma$-finite kernels. They discuss in details the problem of proving the measurability needed to define the composition of kernels, which we omitted in this paper. Both of those formalizations build classes of kernels and some of their compositions (corresponding to what we present in Section 2) but do not discuss disintegration of kernels or conditional distributions.
A small difference between the implementations of both [hirata2023semantic] and [affeldt2025semantics] and ours is that they do not define a kernel as a measurable function into a measurable space of measures, but as a function into measures that satisfies the set-based measurability condition directly (for every measurable set $B$, the function $x\mapsto\kappa(x)(B)$ is measurable). Our definition uses the measurable space on measures to write that the function $x\mapsto\kappa(x)$ is measurable instead of spelling out what it means in terms of evaluation on measurable sets.
In our implementation, only Kernel is a structure, and then s-finite, finite and Markov kernels are properties of Kernel (which are implemented as typeclasses). In contrast, [affeldt2025semantics] define structures for each class of kernels they introduce. We are not knowledgeable enough about the Rocq language to comment meaningfully on the significance of that design choice.
[affeldt2025semantics] mention several composition operations on kernels. Our composition-product is the product of their Theorem 8.1, while our composition corresponds to the operation of their Theorem 8.3. Interestingly, the central operation in their work (Theorem 5.1) is not exactly one we implemented. It takes two s-finite kernels $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}$ and $\eta:\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$ and produces a kernel $\kappa;\eta:\mathcal{X}\rightsquigarrow\mathcal{Z}$, which is such that for all measurable functions $f:\mathcal{Z}\to\mathbb{R}_{+}$, for all $x\in\mathcal{X}$,
$\int_{z}f(z)d(\kappa;\eta)(x)=\int_{y\in\mathcal{Y}}\int_{z\in\mathcal{Z}}f(z)d\eta(x,y)d\kappa(x)\>.$
This is also the main composition operation discussed by [staton2020probabilistic]. We could implement it by taking the projection of the composition-product to the second coordinate ($(\kappa\otimes_{k}\eta)_{\mathrm{snd}}$), but we did not see a need for it in our work. This is probably due to our focus on kernels as a tool for probability theory, while those works are motivated by probabilistic programming languages.
7.2 The Ionescu-Tulcea theorem
We highlight a recent application of the kernel framework: the Ionescu-Tulcea theorem, which was implemented in Lean and added to Mathlib by [marion2025formalization].
Theorem 7.1 (Ionescu-Tulcea).
Let $(X_{n})_{n\in\mathbb{N}}$ be a family of measurable spaces. Let $(\kappa_{n})_{n\in\mathbb{N}}$ be a family of Markov kernels such that for any $n$, $\kappa_{n}$ is a kernel from $\prod_{i=0}^{n}X_{i}$ to $X_{n+1}$. Let $\pi_{[1,n]}:\prod_{i=1}^{\infty}X_{i}\to\prod_{i=1}^{n}X_{i}$ be the projection on the first $n$ coordinates. Then there exists a unique Markov kernel $\xi:X_{0}\rightsquigarrow\prod_{i=1}^{\infty}X_{i}$ such that for any $n\geq 1$, the pushforward of $\xi$ by $\pi_{[1,n]}$ is $\pi_{[1,n]*}\xi=\kappa_{0}\otimes_{k}\ldots\otimes_{k}\kappa_{n-1}$.
In the formalization of that Theorem, an important difficulty is the lack of associativity of products of types, as explained in Section 2.2 about the composition-product. The article [marion2025formalization] discusses this issue in detail and presents a formalization strategy to avoid inserting associators everywhere.
The main application of the Ionescu-Tulcea theorem in Mathlib is to define the product measure of an infinite family of probability measures (Measure.infinitePi), indexed by an arbitrary, possibly uncountable, type.